Factor​ Completely:
Factor the polynomial completely.
5x^3-10x^2-7x+14


Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Factor​ Completely:
Factor the polynomial completely.
5x^3-10x^2-7x+14
=================
Factoring is learned by doing it, not having someone do it for you.
Look for patterns.
---
Notice the 2nd coeff is 2 times the 1st, and the 4th is 2 times the 3rd.
That's a good hint.
============================
Update: another "tutor" gave you the answer.
No need to learn anything.
============================
I would not bet $0.50 that the student now knows how to do this type of problem.
He/she got the answer, moved on.

Answer by ikleyn(52792)

.

  5x^3 - 10x^2 - 7x + 14 = 



    To factor this polynomial, apply the grouping method.

    First step is to group the terms appropriately :



= (5x^3 - 10x^2) - (7x - 14) = 



    Factor out 5x^2 from the first parentheses;  factor out 7 from the second parentheses. You will get



= 5x^2*(x-2) - 7*(x-2) = (x-2)*(5x^2 - 7).      ANSWER


Factoring is completed and the problem is just solved.


/\/\/\/\/\/\/\/

Alan said after my post that you do not need to learn anything, since I just done everything for you.

He is not right: You still need learn something.

Namely, you need learn my solution and the way on how I solved the problem, to be able to solve next similar problem on your own !

Also you need to learn all relevant words and terms that I used on the way.


Happy learning ! ! !


And do not forget to post your "Thanks" to me for my lesson.